[Ncep.nhc.nco_contacts] (no subject)

Mark DeMaria - NOAA Federal mark.demaria at noaa.gov
Fri Jan 8 13:37:14 UTC 2016


Here is today's update on NHC's computer operations.

Mark

-----
TAFB:

The ftp gate processing still does not work on the derecho N-AWIPS
workstation. NCO made some adjustments yesterday, but they did not work. We
will continue working on this.

HSU:

NHC is doing back up for HFO on Monday, including the advisories for
tropical storm Pali. Since we don't have the Himawari data in N-AWIPS yet,
we are extending the GOES-W IR and Vis sectors further west to make sure
they cover TS Pali.

Other Issues:

Work continues on the AC system in the computer room.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160108/3e959124/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list